OpenStudy (anonymous):

Which of the following exponential functions goes through the points (1, 6) and (2, 12)? f(x) = 3(2)x f(x) = 2(3)x f(x) = 3(2)−x f(x) = 2(3)−x

OpenStudy (anonymous):

@Astrophysics

OpenStudy (astrophysics):

Ok now this time, they are all exponents?

OpenStudy (astrophysics):

If they are exponents you should start putting "^" so 3(2)^x

OpenStudy (anonymous):

okay then what

OpenStudy (astrophysics):

So specify your options...

OpenStudy (astrophysics):

\[f(x) = 3(2)^x~~~~or~~~~f(x) = 3(2)x\] put a hat on all the ones that look like the first function because I really don't want to do what happened last time again

OpenStudy (anonymous):

i say A.?

OpenStudy (astrophysics):

What? Ok sorry but I think you need to look up exponents and what they are, good luck.